OpenStudy (emilyisanemoji):

What is the value of B? b/-3=1.5 will give medals and will fan

OpenStudy (photon336):

\[\frac{ b }{ -3 } = \frac{ 3 }{ 2 }\]

OpenStudy (emilyisanemoji):

Help?

OpenStudy (photon336):

@Emilyisanemoji I re-wrote 1.5 as 3/2

OpenStudy (photon336):

the first step is to get b by itself

OpenStudy (emilyisanemoji):

ok

OpenStudy (photon336):

see that -3 in the denominator?

OpenStudy (emilyisanemoji):

do i multiply them both by -3?

OpenStudy (photon336):

yes! good work

OpenStudy (photon336):

\[-\cancel\3*(\frac{ b }{ -\cancel\3 }) = \frac{ 3 }{ 2 }*(\frac{-3 }{ 1 })\]

OpenStudy (emilyisanemoji):

?

OpenStudy (photon336):

\[b = \frac{ 3 }{ 2 }*(-3)\]

OpenStudy (emilyisanemoji):

i dont get it

OpenStudy (photon336):

we multiplied both sides by -3

OpenStudy (emilyisanemoji):

ya

OpenStudy (emilyisanemoji):

its -4.5 right?

OpenStudy (photon336):

yep \[-3*(\frac{ 3 }{ 2 }) = \frac{ -3*3 }{ 2 } = \frac{ -9 }{ 2 }\] \[-4.5 \]

OpenStudy (emilyisanemoji):

YAY!